Write the slope-intercept equation for the line that passes through (5, 4) and (7, 8). Please show all of your work
Answer by mananth(16946)


slope m =(y2-y1)/(x2-x1)
(8-4)/(7-5)
m=2

Plug value of the slope and point (5,4) in
Y =mx+b
4.00=10+ b
b=4-10
b=-6
So the equation will be
Y =2x-6